What is the pH of Acetic Acid has a Molarity of 0.0001
1
Expert's answer
2020-04-09T13:10:55-0400

Acetic acid = CH3COOH


Solution:

Acetic acid (CH3COOH) is a weak acid in aqueous media.

Therefore, CH3COOH ionizes but less than 100% to produce protons (H+).

The acid dissociation constant for CH3COOH is Ka = 1.8×10−5.

Let's set up a reaction table for the following ionization: CH3COOH ↔ H+ + CH3COO-


Then,

x*x = 1.8*10-5 * (0.0001 - x);

x2 + 1.8*10-5*x - 1.8*10-9 = 0

Solving this quadratic equation, we get that x = 0.00003437 = 3.437*10-5

x = [H+] = 3.437*10-5 M.


pH = −log[H+].

Therefore,

pH = −log[H+] = −log[3.437*10-5] = 4.46

pH = 4.46


Answer: pH = 4.46.